Prove that MSE(Ô) = Var() + Bias(0)2, i.e., E[(ôn – 0)21 = E[(ên – E(0))21 + [E(Ô) – 012.
Answer #1

Question 11 + S (6) = E (6 – 8)] = E [6- E[6]+E() –-)] El 16– E (6)}+ {E(6) – 03 +248-E16)}fE (6)-0}] El{6-E (0)²)+ EXE (O)-